OpenStudy (anonymous):

make a substitution to express the integrand as a rational function and then evaluate the integral \[\int_{9}^{16} \frac{\sqrt(x)}{x-4} dx\]

OpenStudy (anonymous):

isn't it already a rational function?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

rational function is not one function over another, it is the ratio of two polynomials

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i guess you should try \(u=\sqrt{x}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

that would make things more complicated i think

OpenStudy (anonymous):

then don't be scared because \(du=\frac{1}{2\sqrt{x}}dx\)

OpenStudy (anonymous):

no it makes it easier

OpenStudy (anonymous):

\[\frac {u}{u^2 -4}\]

OpenStudy (anonymous):

since \(u=\sqrt{x}\) you get \(2\sqrt{x}du=dx\) or \[2u=dx\]

OpenStudy (anonymous):

what about the dx?

OpenStudy (anonymous):

lets go back to the start

OpenStudy (anonymous):

i see it now...

OpenStudy (anonymous):

\[u=\sqrt{x}\] \[du=\frac{1}{2\sqrt{x}}dx\] \[2\sqrt{x}du=dx\] and since \(u=\sqrt{x}\) this becomes \(2udu=dx\)

OpenStudy (anonymous):

\frac {2u^2}{u^2 -4}

OpenStudy (anonymous):

\[\frac {2u^2}{u^2 -4} \]

OpenStudy (anonymous):

so your integral is \[\int\frac{u^2du}{u^2-4}\]

OpenStudy (anonymous):

oops i dropped the 2 you are right don't forget to change the limits of integration

OpenStudy (anonymous):

plug in 16 and 9 to sqrtx ?

OpenStudy (anonymous):

yes, exactly doesn't what work out nicely?

OpenStudy (anonymous):

yep...thanks again

OpenStudy (anonymous):

yw integral should be ok from there right? start with \[2\int_3^4\frac{u^2du}{u^2-4}\] and divide

OpenStudy (anonymous):

actually there is still some annoying pfd to do , but it is doable